OpenStudy (anonymous):

Solve using law of sines A = 41°, B =61°, c = 19

OpenStudy (anonymous):

law of sines \[\frac{ \sin A }{ a }=\frac{ \sin B }{ b }=\frac{ \sin C }{ c }\] \[\frac{ \sin 41 }{ a }=\frac{ \sin 61 }{ b }=\frac{ \sin C }{ 19 }\]

OpenStudy (anonymous):

\[ A+B+C=180^\circ\\ 41^\circ+61^\circ+C=180^\circ\\ C=180^\circ-41^\circ-61^\circ\\ \ \ \ =78^\circ\\ \frac{a}{\sin A}=\frac{b}{\sin B}=\frac{c}{\sin C}\\ \\a=c\frac{\sin A}{\sin C}\\ \\b=c\frac{\sin B}{\sin C}\\ \text{The rest is left as exercise to you.} \]

OpenStudy (anonymous):

nvm i think i got the answer